# Stability of non-supersymmetric vacua from calibrations

### v2

@@ 5.2.2 The solutions
Branch 2 consists of the supersymmetric and skew-whiffed solutions (with the minus and plus sign, respectively). Branch 1 breaks supersymmetry with a stretched SE; it was found in a slightly different language in [26, Sec. 3]. Notice that the case $s_{1}=1$ has $G_{3}=F_{3}-\mathrm{i}\mathrm{e}^{-\phi}H=\tilde{f}_{3}\omega$.
@@ 5.3.2 The solutions
Due to (5.31d), there are again two branches:
$\begin{array}[]{l}f_{31}=f_{32}=h_{1}=h_{2}=0\,,\qquad\kappa_{1}=\kappa_{2}=1\,,\\[5.0pt] \sigma^{2}=\frac{2\alpha^{2}}{p_{1}^{2}\alpha^{4}+2p_{2}^{2}}\,,\qquad f_{5}^{2}=\frac{4(p_{1}^{2}\alpha^{4}+p_{2}^{2})}{g_{s}^{2}R^{2}\alpha^{2}(p_{1}^{2}\alpha^{4}+2p_{2}^{2})}\,,\\[10.0pt] p_{1}^{2}\alpha^{4}(\alpha^{2}-2)+p_{2}^{2}(2\alpha^{2}-1)=0\end{array}\quad(\text{Branch }1)$ (5.32)
and
$\begin{array}[]{l}p_{2}f_{31}=-p_{1}f_{32}\alpha^{2}\,,\qquad p_{2}h_{1}=-p_{1}h_{2}\alpha^{2}\,,\qquad f_{5}=0\,,\\[5.0pt] h_{2}=g_{s}f_{32}\,\qquad f_{32}^{2}=\frac{2p_{2}^{2}\sigma^{2}}{3g_{s}^{2}R^{2}\alpha^{4}}\,,\qquad\sigma^{2}=\frac{6\kappa_{2}\alpha^{2}}{-p_{1}^{2}\alpha^{4}+6p_{2}^{2}}\,,\\[10.0pt] p_{1}^{2}\alpha^{4}(\kappa_{1}\alpha^{2}+6\kappa_{2})-p_{2}^{2}(6\kappa_{1}\alpha^{2}+\kappa_{2})=0\,.\end{array}\quad(\text{Branch }2)$ (5.33)
In both branches, $\alpha$ is determined implicitly by the last equation.
The solutions in Branch 1 were found by Romans [26, Sec. 2]: they are of Freund–Rubin type, with $M_{5}$ Einstein (but not Sasaki). Those in Branch 2 appear to be new. Imposing positivity and reality of the parameters shows that $(\kappa_{1},\kappa_{2})=(1,1)$, $(0,1)$, (1,1)(-1,1) are allowed (as well as the symmetric $(1,0)$, (1,1)(1,-1)). [^25] In deriving (5.33) we have assumed $\kappa_{2}\neq 0$; the case $\kappa_{2}=0$ can be studied separately, and is related by (5.28) to the case $\kappa_{1}=0$ .
For Branch 1, flux quantization gives
@@
where $\chi_{a}=\chi(\Sigma_{a})=(2-2g_{a})$. For Branch 2, we need to choose a potential for $H$ in (5.29); we can take
$B=hR^{3}(-e^{5}\wedge(p_{1}a_{1}-p_{2}a_{2})-p_{1}p_{2}\sigma a_{1}\wedge a_{2})+b(-q_{2}j_{1}+q_{1}j_{2})\,;$ (5.35)
[…]
$\frac{\Delta y}{2\pi}R^{3}\sigma f_{3}p_{a}\chi_{a}\in\mathbb{Z}\ \ \text{(no sum)}\,,\qquad\frac{\Delta y}{2\pi^{2}}p_{1}p_{2}bf_{3}\sigma R^{5}\in\mathbb{Z}\,.$ (5.36)
@@ 5.3.3 Stability
We take $\psi^{2}_{+}=\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{-\mathrm{i}J}$, with $J=R^{2}(j_{1}\pm\alpha^{2}j_{2})$.
First we briefly consider Branch 1. D3 stability was considered in (5.7). Note that $f_{5}$ is the only non-zero RR field, so $r_{\mathrm{D3}}=1$ and the D3 is only marginally stable. This is of course at danger of becoming unstable once string corrections are taken into account. These solutions were anyway found to be perturbatively unstable [74]. In view of all this, we are not going to give details for D5-brane stability; our conclusion is that it does in fact hold.
We focus instead on Branch 2. Here D3 stability is very solid, as $f_{5}=0$, so $r_{\mathrm{D3}}=0$. For D5 bound states, we take $2\pi f_{\mathrm{ws}}=f(-q_{2}j_{1}+q_{1}j_{2})$, similar to the closed term in (5.35); for simpler expressions we take $b=0$. For the generic case $p_{a}\neq 0$, again by the Gysin sequence (5.12) there is only one two-cycle $B_{2}\subset M_{5}$; its Poincaré dual is proportional to $\eta\wedge(p_{1}j_{1}-p_{2}j_{2})$, which is closed but not exact (again if $p_{a}\neq 0$). Now
$\int_{B_{2}}\frac{4}{L}\mathrm{e}^{4A-\phi}\mathrm{e}^{-{\mathcal{F}}}\wedge\mathrm{Re}\psi^{2}_{+}$ $=\sqrt{f_{31}^{2}+f_{32}^{2}}\int_{B_{2}}(-\cos\theta{\mathcal{F}}+\sin\theta J)$ (5.37)
$=\sqrt{f_{31}^{2}+f_{32}^{2}}(-\cos\theta fp_{a}q_{a}+\sin\theta R^{2}(-p_{2}\pm p_{1}\alpha^{2}))\int_{M_{5}}\eta\wedge j_{1}\wedge j_{2}\,.$
@@
$\int_{B_{2}}(\mathrm{e}^{-\mathcal{F}}\wedge*\lambda F)=-R^{2}\int_{B_{2}}(\alpha^{2}f_{31}j_{2}+f_{32}j_{1})=-R^{2}(p_{1}f_{31}^{2}-p_{2}f_{32})\int_{M_{5}}\eta\wedge j_{1}\wedge j_{2}\,.$ (5.38)
After some simplification we find
$r_{\mathrm{D5}}=\frac{R^{2}\sqrt{p_{1}^{2}\alpha^{4}+p_{2}^{2}}}{\sqrt{(p_{a}q_{a})^{2}f^{2}+R^{4}(\mp p_{1}\alpha^{2}+p_{2})^{2}}}\,.$ (5.39)
This is <1<1 if we choose $\mp=\mathrm{sign}(p_{1})$.
@@ 6.2 A Minkowski4 application
[31, Sec. 2.3] parameterizes the piecewise-linear function as

### v3

@@ 5.2.2 The solutions
Branch 2 consists of the supersymmetric and skew-whiffed solutions (with the minus and plus sign, respectively). Branch 1 breaks supersymmetry with a stretched SE; it was found in a slightly different language in [59, Sec. 3]. Notice that the case $s_{1}=1$ has $G_{3}=F_{3}-\mathrm{i}\mathrm{e}^{-\phi}H=\tilde{f}_{3}\omega$.
@@ 5.3.2 The solutions
We obtain the solution
$\begin{array}[]{l}f_{31}=f_{32}=h_{1}=h_{2}=0\,,\qquad\kappa_{1}=\kappa_{2}=1\,,\\[5.0pt] \sigma^{2}=\frac{2\alpha^{2}}{p_{1}^{2}\alpha^{4}+2p_{2}^{2}}\,,\qquad f_{5}^{2}=\frac{4(p_{1}^{2}\alpha^{4}+p_{2}^{2})}{g_{s}^{2}R^{2}\alpha^{2}(p_{1}^{2}\alpha^{4}+2p_{2}^{2})}\,,\\[10.0pt] p_{1}^{2}\alpha^{4}(\alpha^{2}-2)+p_{2}^{2}(2\alpha^{2}-1)=0\,.\end{array}$ (5.32)
$\alpha$ is determined implicitly by the last equation. These were found by Romans [59, Sec. 2]: they are of Freund–Rubin type, with $M_{5}$ Einstein (but not Sasaki).
Flux quantization gives
@@
where $\chi_{a}=\chi(\Sigma_{a})=(2-2g_{a})$.
@@ 5.3.3 Stability
These solutions were found to be perturbatively unstable [34], so our bubble stability analysis will be brief.
We take $\psi^{2}_{+}=\mathrm{e}^{\mathrm{i}\theta}\mathrm{e}^{-\mathrm{i}J}$, with $J=R^{2}(j_{1}\pm\alpha^{2}j_{2})$. D3 stability was considered in (5.7). Note that $f_{5}$ is the only non-zero RR field, so $r_{\mathrm{D3}}=1$ and the D3 is only marginally stable. This is of course at danger of becoming unstable once string corrections are taken into account.
For D5 bound states, we take $2\pi f_{\mathrm{ws}}=f(-q_{2}j_{1}+q_{1}j_{2})$. For the generic case $p_{a}\neq 0$, again by the Gysin sequence (5.12) there is only one two-cycle $B_{2}\subset M_{5}$; its Poincaré dual is proportional to $\eta\wedge(p_{1}j_{1}-p_{2}j_{2})$, which is closed but not exact. Now
$\int_{B_{2}}\frac{4}{L}\mathrm{e}^{4A-\phi}\mathrm{e}^{-{\mathcal{F}}}\wedge\mathrm{Re}\psi^{2}_{+}$ $=|f_{5}|\int_{B_{2}}(-\cos\theta{\mathcal{F}}+\sin\theta J)$ (5.34)
$=|f_{5}|(-\cos\theta fp_{a}q_{a}+\sin\theta R^{2}(-p_{2}\pm p_{1}\alpha^{2}))\int_{M_{5}}\eta\wedge j_{1}\wedge j_{2}\,.$
@@
$\int_{B_{2}}\mathrm{e}^{-\mathcal{F}}\wedge*\lambda F=-ff_{5}\int_{B_{2}}(-q_{2}j_{1}+q_{1}j_{2})=-ff_{5}p_{a}q_{a}\int_{M_{5}}\eta\wedge j_{1}\wedge j_{2}\,.$ (5.35)
Maximizing in $\theta$, we find
$r_{\mathrm{D5}}=\frac{2fp_{a}q_{a}}{\sqrt{(p_{a}q_{a})^{2}f^{2}+R^{4}(\mp p_{1}\alpha^{2}+p_{2})^{2}}}\,.$ (5.36)
which is always <1<1.
@@ 6.2 A Minkowski4 application
[46, Sec. 2.3] parameterizes the piecewise-linear function as